OpenStudy (anonymous):

A train on a straight, level track has an initial speed of 35 km/h. a uniform acceleration of 1.5m/s^2 is applied while the train travels 200 m. A)what is the speed of the train at the end of this distance? B)how long did it take the train to travel 200m? The answers are A) 26.3m/s B) 11.1s. can someone help me understand how they got these answers please. Thanks

OpenStudy (anonymous):

(a)v^2 = u^2 + 2as u=35km/h = 35000m/h = 583.33m/min = 9.72m/s a=1.5ms^-2 s=200m v=26.4m/s (b)v=u+at 26.4=9.72+1.5t t=11.12s

OpenStudy (anonymous):

Thanks that helps alot

OpenStudy (anonymous):

U are welcome